OpenStudy (anonymous):

divide 2/3 and 1/2 write your answer in lowest terms

OpenStudy (anonymous):

4/3

OpenStudy (anonymous):

yeah but i says write your answer as a mixed or whole number

OpenStudy (anonymous):

do you know how to write \(\frac{4}{3}\) as a mixed number?

OpenStudy (anonymous):

you divide 3 into 4 right?

OpenStudy (anonymous):

yes, and take the remainder and put it over 3

OpenStudy (shamil98):

(2/3)/(1/2) = (2/3)(2/1) when dividing fractions you multiply by the reciprocal. so then it would be 4/3.

OpenStudy (anonymous):

1/3 right

OpenStudy (shamil98):

the numerator is the remainder if it is higher than the denominator.

OpenStudy (shamil98):

So, you would write it as 1 1/3

OpenStudy (anonymous):

so how would i change it to a mixed or whole nuber

OpenStudy (anonymous):

thanks

OpenStudy (shamil98):

how many times does 3 go into 4? once right? So then you have your whole number on the side which is 1

OpenStudy (anonymous):

oh i get it lool

OpenStudy (shamil98):

then you right the remainder as the numerator and 3 as the denominator. and you're welcome.